OpenStudy (anonymous):

CHECK MY ANSWER y=3sin(x-(pi/4)) find the zeros 1/3+/- npi/4

OpenStudy (anonymous):

it is zero when \(x-\frac{\pi}{4}=n\pi\)

OpenStudy (anonymous):

so x=npi+pi/4

OpenStudy (jdoe0001):

notice the shift